### v1

# Learning More Effective Representations for Dense Retrieval through Deliberate Thinking Before Search

Abstract
Recent dense retrievers usually thrive on the emergency capabilities of Large Language Models (LLMs), using them to encode queries and documents into an embedding space for retrieval. These LLM-based dense retrievers have shown promising performance across various retrieval scenarios. However, relying on a single embedding to represent documents proves less effective in capturing different perspectives of documents for matching. In this paper, we propose Deliberate Thinking based Dense Retriever (Debater), which enhances these LLM-based retrievers by enabling them to learn more effective document representations through a step-by-step thinking process. Debater introduces the Chain-of-Deliberation mechanism to iteratively optimize document representations using a continuous chain of thought. To consolidate information from various thinking steps, Debater also incorporates the Self Distillation mechanism, which identifies the most informative thinking steps and integrates them into a unified text embedding. Experimental results show that Debater significantly outperforms existing methods across several retrieval benchmarks, demonstrating superior accuracy and robustness. All codes are available at https://github.com/OpenBMB/DEBATER.
Learning More Effective Representations for Dense Retrieval through Deliberate Thinking Before Search
Yifan Ji1thanks:   indicates equal contribution., Zhipeng Xu [^1] footnotemark: 1, Zhenghao Liu1thanks:   indicates corresponding author., Yukun Yan2, Shi Yu2, Yishan Li3, Zhiyuan Liu2, Yu Gu1, Ge Yu1, Maosong Sun2 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3ModelBest.Inc
1 Introduction
Dense retrieval models encode both queries and documents into a dense embedding space and measure their similarity to retrieve relevant documents Karpukhin et al. (2020); Zhao et al. (2024); Xiong et al. (2021), demonstrating strong effectiveness in various downstream NLP tasks, such as open-domain question answering Chen and Yih (2020), fact verification Liu et al. (2020), and web search Chen et al. (2024). However, recent findings have shown that dense retrievers suffer from significant performance degradation when applied to new tasks or domains Su et al. (2023), raising concerns about their versatility Luo et al. (2024); Khramtsova et al. (2024).
Figure 1: The Illustration of Our Deliberate Thinking based Dense Retriever (Debater). Debater leverages the reasoning capability of LLM to conduct fine-grained document representations for retrieval.
Large Language Models (LLMs), such as ChatGPT Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated extraordinary emergent capabilities Wei et al. (2022a); Zhao et al. (2023), inspiring researchers to leverage them to enhance the task and domain generalization of dense retrievers Zhu et al. (2023); Khramtsova et al. (2024). In particular, existing work has focused on prompting LLMs to generate dense representations for retrieval Zhuang et al. (2024). These methods typically use task-specific instructions or in-context demonstrations to guide LLMs in generating task- and domain-aware embeddings. To learn more tailored representations for dense retrieval, researchers further focus on optimizing LLM-based retrievers using relevance labels Ma et al. (2024); Neelakantan et al. (2022); Li et al. (2025). These methods exploit the superior reasoning abilities of LLMs, achieving impressive performance across various retrieval tasks Wang et al. (2024a); Zhu et al. (2023); Luo et al. (2024). Recent studies suggest that LLMs pose strong reasoning capability, particularly implemented by their step-by-step thinking Kudo et al. (2024); Wei et al. (2022b). LLM-based retrievers typically rely on the hidden state of the end-of-sequence token as both query and document representations. Nevertheless, only relying on one embedding usually shows less effectiveness in representing documents from different views that can match queries Zhang et al. (2022); Khattab and Zaharia (2020).
In this paper, we propose a Deliberate Thinking based Dense Retriever (Debater) model to learn more effective document representations through deliberately thinking step-by-step before retrieval. As shown in Figure 1, our method stimulates LLMs to conduct the reasoning process, enabling them to generate more fine-grained document representations for retrieval. Specifically, Debater introduces the Chain-of-Deliberation mechanism to encourage LLMs to conduct deliberate thinking by autograssively decoding the document representations. Then Debater utilizes the Self Distillation mechanisms to gather all information from previous steps and compress them into the document embedding at the last step.
Our experiments show that Debater achieves comparable or even better retrieval performance than the baseline methods implemented by larger-scale LLMs, highlighting its effectiveness. Our further analyses show that both Chain-of-Deliberation and Self Distillation play important roles in Debater and appropriately increasing the thinking steps can benefit these LLM-based dense retrieval models. Thriving on autograssively decoding different document representations during thinking, the document representations can be gradually refined to be more effective. By incorporating our Self Distillation, LLMs show the ability to capture different key information at different thinking steps and gather all crucial semantics from different steps to the final document representations.
2 Related Work
Dense retrieval Karpukhin et al. (2020); Xiong et al. (2021); Su et al. (2023) has proven effective in various NLP downstream tasks Liu et al. (2020); Chen et al. (2024); Guu et al. (2020). However, the versatility of dense retrievers remains a challenge that hinders their progress Luo et al. (2024); Lee et al. (2024), particularly their inability to generate task- and domain-specific embeddings and return suitable results Su et al. (2023); Luo et al. (2024); Tao et al. (2024). To address this limitation, prior work has focused on conducting fine-grained data curation to fine-tune dense retrievers with multi-task instructions Su et al. (2023); Asai et al. (2023). However, obtaining high-quality relevance labels can be difficult for training dense retrievers Yu et al. (2022); Gao et al. (2023); Wang et al. (2024a).
Recent research has shifted towards using LLMs as the backbone for dense retrievers Tao et al. (2024), thriving on their strong emergence capabilities. Some studies attempt to directly prompt LLMs to generate embeddings for retrieval Zhuang et al. (2024). However, prompt-based approaches cannot leverage pre-existing retrieval signals, limiting their effectiveness Zhu et al. (2023). In contrast, recent efforts have focused on fine-tuning LLMs for dense retrieval tasks Wang et al. (2024a); Ma et al. (2024); Li et al. (2024), or designing additional pretraining tasks to transform LLMs into dense retrievers BehnamGhader et al. (2024), achieving strong retrieval performance and generalization capabilities. However, existing methods typically extract the last hidden state of the end-of-sequence token as the dense representation Ma et al. (2024); Luo et al. (2024), which is not always effective for fully representing documents from different perspectives to match queries Zhang et al. (2022); Khattab and Zaharia (2020). The exploration of different document representations, such as leveraging the reasoning ability of LLMs, remains an underexplored area.
Figure 2: The Architecture of Deliberate Thinking based Dense Retriever (Debater). Debater uses both the Chain-of-Deliberation (CoD) and Self Distillation (SD) for training LLMs.
To enhance the reasoning capability of LLMs, one approach is to generate intermediate reasoning steps using Chain-of-Thought (CoT) Wei et al. (2022b) or its variants Chen et al. (2023); Zhang et al. (2024). CoT allows LLMs to delay final answers by engaging in reasoning Kudo et al. (2024), improving response accuracy Wei et al. (2022b); Chu et al. (2024). However, these approaches operate within the language space and often require generating tens or even hundreds of additional tokens, which can hinder their ability to meet the latency requirements of dense retrievers. Current research is exploring the integration of CoT reasoning into a continuous latent space Hao et al. (2024); Xie et al. (2024) to enhance computational efficiency. Building on these advancements, our Debater focuses on latent reasoning chains, encouraging LLM-based retrievers to think step-by-step to enhance the dense representations of documents.
3 Methodology
In this section, we introduce our Deliberate Thinking based Dense Retriever (Debater). We first introduce the preliminary of LLM-based dense retrieval (Sec. 3.1). Then we describe our deliberation thinking based embedding learning method used by Debater (Sec. 3.2).
3.1 Preliminary of Dense Retrieval with Large Language Models as Foundations
Given a query $q$ and a document collection $\mathcal{D}$, the goal of the retrieval task is to identify a subset of documents most relevant to the query.
LLM-based dense retrievers typically map both the query $q$ and document $d$ into a shared latent space for retrieval, where the query embedding $h^{q}$ and document embedding $h^{d}$ are defined as:
[…]
The ranking score $f(q,d)$ between the query embedding $h^{q}$ and the document embedding $h^{d}$ is calculated as:
$f(q,d)=sim(h^{q},h^{d}),$ (2)
where sim denotes the similarity function. In Debater, we use cosine similarity to measure the similarity between queries and documents, which is also employed in previous works Wang et al. (2024a); BehnamGhader et al. (2024). Subsequently, we train the LLMs contrastively to maximize the probability of retrieving the positive document $d^{+}$ over the negative document $d^{-}$:
$p(d^{+}|q,{d^{+}}\cup\mathcal{D}^{-})=\frac{e^{f(q,d^{+})}}{e^{f(q,d^{+})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f(q,d^{-})}},$ (3)
where $\mathcal{D}^{-}$ denotes the set of negative documents, typically obtained via in-batch negative sampling Karpukhin et al. (2020).
Current LLM-based dense retrievers typically use the last hidden state corresponding to the end-of-sequence token (</s>) as the dense representation. However, they do not fully exploit the reasoning capabilities of LLMs, which helps conduct more effective representations by learning information from diverse views of documents.
3.2 Enhancing Dense Retriever through Deliberate Thinking
In this subsection, we introduce the Deliberate Thinking based Dense Retriever (Debater), which aims to unleash the reasoning ability of LLMs and generate more fine-grained document representations. As shown in Figure 2, Debater consists of two modules to enhance the LLM-based dense retriever: Chain-of-Deliberation (CoD) and Self Distillation (SD).
Chain-of-Deliberation. To enhance these LLM-based dense retrievers, Debater introduces the Chain-of-Deliberation (CoD) approach, which delays the computation of document embeddings by performing several steps of reasoning.
Specifically, CoD incorporates a sequence of prompt tokens $\{t_{1},t_{2},\dots,t_{m}\}$ to stimulate the reasoning capability of LLMs when representing the document $d$. These tokens $\{t_{1},t_{2},\dots,t_{m-1}\}$ serve as intermediate thinking steps, encouraging the model to think step-by-step before producing the final document embedding at the $m$-th step:
$h_{m}^{d}=\text{LLM}(X,t_{1},t_{2},\dots,t_{m-1},t_{m}),$ (4)
where $m$ is a hyperparameter to control the thinking depth. An appropriate choice of $m$ is crucial to avoid overthinking or under-optimization.
During training, we first calculate the similarity score between query representation $h^{q}$ and the document representation $h^{d}_{i}$ at the $i$-th thinking step:
$f(q,d(t_{i}))=sim(h^{q},h^{d}_{i}).$ (5)
Next, we gather all similarity scores using the decoded document representations $\{h^{d}_{1},...,h^{d}_{m}\}$. We then select the most useful thinking step from CoD and use the corresponding embedding as the document representation to compute the training loss. The relevance scores $f_{\text{max}}(q,d)$ between the query and the document are computed as:
$f_{\text{max}}(q,d)=\max_{1\leq i\leq m}sim(h^{q},h^{d}_{i}),$ (6)
The LLM is optimized by minimizing the contrastive training loss:
$\mathcal{L}_{c}=-\text{log}\frac{e^{f_{\text{max}}(q,{d^{+}})}}{e^{f_{\text{max}}(q,{d^{+}})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f_{\text{max}}(q,{d^{-}})}}.$ (7)
Self Distillation. Although the final token of the Chain-of-Deliberation aggregates information from all thinking steps through autoregressive decoding, it may overlook crucial reasoning cues presented in embeddings decoded at earlier steps.
To address this, we introduce Self Distillation (SD), a strategy for distilling knowledge from different thinking steps into the final document representation $h^{d}_{m}$. Specifically, we use the most informative thinking step as the teacher to guide the representation learning of the final token in CoD, thereby enhancing the document representation.
For the query $q$, we compute the ranking probability of the $i$-th document $d_{i}$ in the document collection $\tilde{\mathcal{D}}=\{d^{+}\}\cup\mathcal{D}^{-}$ as:
$P(d_{i}|q)=\frac{e^{f_{\text{max}}(q,d_{i})}}{\sum_{d_{j}\in\tilde{\mathcal{D}}}e^{f_{\text{max}}(q,d_{j})}},$ (8)
where $|\tilde{\mathcal{D}}|=k$. This yields a probability distribution $P(\tilde{\mathcal{D}}|q)$ over the $k$ documents:
$P(\tilde{\mathcal{D}}|q)=\left[P(d_{1}|q),P(d_{2}|q),\dots,P(d_{k}|q)\right].$ (9)
Each value $P(d_{i}|q)$ represents the ranking probability of the $i$-th document $d_{i}$ using the document representations from all thinking steps $\{h^{d}_{1},...,h^{d}_{m}\}$ of CoD that yield a higher similarity with the query. Concurrently, we compute the rank probability of $d_{i}$ using the last-token embedding $h^{d}_{m}$ from CoD:
$Q(d_{i}(t_{m})|q)=\frac{e^{f(q,d_{i}(t_{m}))}}{\sum_{d_{j}\in\{d^{+}\}\cup\mathcal{D}^{-}}e^{f(q,d_{j}(t_{m}))}}.$ (10)
Then we can obtain the ranking probability distribution $Q(\tilde{\mathcal{D}}|q)$ as well:
$Q(\tilde{\mathcal{D}}|q)=\left[Q(d_{1}|q),Q(d_{2}|q),\dots,Q(d_{k}|q)\right].$ (11)
We then minimize the Kullback-Leibler (KL) divergence between two probability distributions $P(\tilde{\mathcal{D}}|q)$ and $Q(\tilde{\mathcal{D}}|q)$:
$\mathcal{L}_{t}={P(\tilde{\mathcal{D}}|q)\cdot\log\frac{P(\tilde{\mathcal{D}}|q)}{Q(\tilde{\mathcal{D}}|q)}},$ (12)
where the Self Distillation loss $\mathcal{L}_{t}$ optimizes the document representation $h_{m}^{d}$ by capturing more crucial matching signals from all thinking steps.
Training. Finally, we train our Debater models by minimizing the following loss $\mathcal{L}$:
$\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{t},$ (13)
where $\mathcal{L}_{c}$ optimizes the CoD, and $\mathcal{L}_{t}$ is used to distill crucial information from the thinking steps into the final dense representation of the document. This combined loss allows Debater to leverage both thinking depth and self-knowledge distillation to improve retrieval performance.
4 Experimental Methodology
In this section, we describe the datasets, baselines, and implementation details for our experiments.
Method ($\rightarrow$) BM25 GTR SGPT PromptReps RepLLaMA Emb-V3 E5-Mistral Debater
Model Size ($\rightarrow$) / 4.8B 5.8B 8B 7B 7B 7B 2.4B 4B
TREC-COVID 0.656 0.501 0.873 0.693 0.847 0.794 0.708 0.795 0.836
NFCorpus 0.325 0.342 0.362 0.330 0.378 0.336 0.353 0.378 0.399
NQ 0.329 0.568 0.524 0.431 0.624 0.580 0.482 0.560 0.561
HotpotQA 0.603 0.599 0.593 0.471 0.685 0.668 0.756 0.678 0.678
FiQA 0.236 0.467 0.372 0.324 0.458 0.388 0.545 0.434 0.462
ArguAna 0.414 0.540 0.514 0.330 0.486 0.508 0.625 0.567 0.562
Touché-2020 0.367 0.256 0.254 0.218 0.305 0.319 0.191 0.211 0.250
Quora 0.789 0.892 0.846 0.805 0.868 0.881 0.895 0.886 0.886
DBPedia 0.313 0.408 0.399 0.377 0.437 0.410 0.477 0.430 0.432
SCIDOCS 0.158 0.161 0.197 0.176 0.181 0.181 0.190 0.197 0.212
FEVER 0.753 0.740 0.783 0.711 0.834 0.876 0.731 0.859 0.857
Climate-FEVER 0.213 0.267 0.305 0.214 0.310 0.289 0.252 0.303 0.294
SciFact 0.665 0.662 0.747 0.657 0.756 0.667 0.744 0.735 0.743
CQADupStack 0.299 0.399 0.381 / / 0.389 / 0.431 0.428
Avg CPT sub 0.485 0.516 0.550 0.466 0.579 0.558 0.571 0.571 0.582
Avg 0.437 0.486 0.511 / / 0.520 / 0.533 0.543
Table 1: Overall Retrieval Performances on BEIR Benchmark. indicates the 11 most representative BEIR tasks used in CPT Neelakantan et al. (2022) and Avg CPT Sub reflects the average performance across these tasks.
Datasets. We train all Debater models using the public portion of the E5 dataset Wang et al. (2024a); Springer et al. (2025), which comprises approximately 1.5M samples. The retrieval effectiveness of Debater is evaluated on the BEIR benchmark Thakur et al. (2021), which includes 18 datasets that span a variety of domains. Our evaluation focuses on the 14 publicly available datasets used for the zero-shot retrieval task. Additional details about the training dataset and evaluation benchmark can be found in Appendix 6.
Evaluation Metrics. To evaluate the retrieval effectiveness of Debater, we use nDCG@10, the standard metric for the BEIR benchmark. The metric implementation follows the pytrec-eval toolkit Gysel and de Rijke (2018), which is consistent with prior work Zhu et al. (2023).
Baselines. We compare Debater with several baseline retrievers implemented with different language models. GTR Ni et al. (2022) employs large dual encoder-only models to build a dense retriever, while SGPT Muennighoff (2022) trains dense retrieval models using decoder-only architectures. Emb-V3111https://cohere.com/blog/introducing-embed-v3 is a commercial text retrieval model provided by Cohere. PromptReps Zhuang et al. (2024) directly prompts LLMs to generate dense representations without supervision. RepLLaMA Ma et al. (2024) and E5-Mistral Wang et al. (2024b) fine-tune LLMs as dense retrievers, using the hidden state of an additional end-of-sequence token to represent the input context. Notably, E5-Mistral is trained on the same dataset as Debater but leverages a larger foundational model.
Implementation Details. We initialize the Debater models with MiniCPM-2.4B and MiniCPM-4B Hu et al. (2024). All Debater models are trained for 1,000 steps using the AdamW optimizer with a batch size of 256. The learning rate follows a cosine decay schedule, with a warm-up phase covering the first 3% of the total iterations, peaking at 2e-4. We train Debater using hybrid negatives, including one hard negative from the E5 dataset and seven in-batch negatives. The CoD length for all Debater models is set to 8. Debater is implemented using the OpenMatch toolkit Yu et al. (2023), with flash-attention Dao et al. (2022) and LoRA Hu et al. (2022) enabled to mitigate memory constraints and improve computational efficiency.
5 Evaluation Results
In this section, we first evaluate the retrieval effectiveness of Debater and then conduct ablation studies to show the roles of different modules in Debater. Then we analyze the characteristics of learned embeddings during thinking step by step.
Method MiniCPM-2.4B MiniCPM-4B
Vanilla w/ SD w/ CoD Debater Vanilla w/ SD w/ CoD Debater
TREC-COVID 0.728 0.822 0.805 0.795 0.747 0.742 0.791 0.836
NFCorpus 0.368 0.368 0.371 0.378 0.379 0.388 0.378 0.399
NQ 0.545 0.531 0.568 0.560 0.533 0.544 0.508 0.561
HotpotQA 0.670 0.656 0.669 0.678 0.564 0.597 0.631 0.678
FiQA-2018 0.406 0.409 0.430 0.434 0.428 0.428 0.413 0.462
ArguAna 0.561 0.526 0.547 0.560 0.569 0.575 0.497 0.562
Touché-2020 0.202 0.250 0.219 0.211 0.195 0.208 0.237 0.250
Quora 0.880 0.788 0.882 0.886 0.886 0.890 0.883 0.886
SCIDOCS 0.191 0.194 0.195 0.197 0.210 0.214 0.198 0.212
Climate-FEVER 0.277 0.310 0.258 0.303 0.211 0.189 0.184 0.294
SciFact 0.715 0.720 0.733 0.735 0.731 0.737 0.730 0.743
Avg 0.504 0.507 0.516 0.522†§ 0.496 0.501 0.495 0.535†‡§
[…]
5.1 Overall Performance
The overall performance of Debater and the baseline retrievers is shown in Table 1. Additional experimental results comparing Debater with other baseline retrievers are provided in Appendix A.4.
Overall, Debater outperforms all baseline retrievers in terms of average retrieval accuracy on BEIR, achieving more than a 2% improvement. This highlights its effectiveness in enhancing the representation capability of LLMs for retrieval. Compared to the prompt-based method PromptReps, these fine-tuned LLM-based methods consistently show improvements, indicating that LLMs also benefit from supervised training to learn more tailored embeddings for retrieval. When compared to E5-Mistral-7B, which is trained on the same E5 corpus as Debater, Debater significantly improves retrieval performance on Trec-COVID, NQ, and FEVER, demonstrating its capability across diverse question-answering scenarios. Notably, when implemented with MiniCPM-2.4B, Debater achieves retrieval performance comparable to that of larger 7B-scale LLM-based dense retrievers while utilizing only 35% of the parameters. This demonstrates that Debater can enhance the representation learning capabilities of smaller-scale LLMs, rather than relying on larger foundational LLMs. Furthermore, when implemented with MiniCPM-4B, the retrieval performance of Debater is improved by 1%, demonstrating that larger models effectively enhance the retrieval capabilities of Debater.
5.2 Ablation Study
As shown in Table 2, we conduct ablation studies to further investigate the roles of Chain-of-Deliberation (CoD) and Self Distillation (SD) modules in Debater.
We compare our Debater with three variations, using MiniCPM-2.4B and MiniCPM-4B as the foundations for building dense retrievers. Both vanilla LLM and MiniCPM w/ CoD models represent documents using the hidden state of the last token and train query and document representations using contrastive training. The key difference between them lies in that MiniCPM w/ CoD performs additional CoD steps before obtaining the document representation. Besides, MiniCPM w/ SD is identical to Debater but removes the CoD steps when generating the document representation.
Compared to vanilla LLM, MiniCPM w/ SD shows almost identical retrieval performance, indicating that relying solely on a few last tokens in the input sequence does not effectively enhance the document representations. This suggests that the special tokens used in CoD serve as prompts that stimulate LLMs to produce more meaningful embeddings. On the other hand, MiniCPM w/ CoD still yields a limited improvement over the vanilla LLM, demonstrating that directly incorporating CoD in representing documents fails to enhance the representation ability of LLMs. After incorporating the Self Distillation mechanism, MiniCPM w/ CoD achieves further improvements, demonstrating its importance in capturing semantics from the different deliberative steps of CoD to optimize the last token as the document representation. Additionally, when using contrastive training to optimize LLMs, the 4B-scale retrieval model performs worse than the 2.4B-scale model. Notably, Debater not only mitigates this performance degradation but also leads to an additional 1.3% improvement, highlighting the effectiveness and robustness of Debater.
5.3 Effectiveness of Chain-of-Deliberation with Different Thinking Depths
In this subsection, we explore how thinking depth affects the effectiveness of Debater. Specifically, we vary the length of the Chain-of-Deliberation (CoD) to train several Debater-2.4B and Debater-4B models and evaluate their retrieval performance on TREC-COVID and FiQA.
(a) TREC-COVID.
(b) FiQA.
(c) NFCorpus.
(d) SCIDOCS.
Figure 3: Retrieval Performance of Debater with Different Thinking Depths. We set the length of the CoD to train different Debater models and evaluate them on different subsets of BEIR.
As illustrated in Figure 3, both Debater-2.4B and Debater-4B exhibit significant and consistent improvements in retrieval performance as the thinking depth increases to 4. This indicates that an appropriate thinking depth effectively activates the reasoning capabilities of LLM-based retrievers, enabling them to generate finer-grained representations of documents. When the thinking depth is further extended to 8, Debater-2.4B reaches a plateau, indicating that it may be nearing its capacity to process more complex or prolonged deliberations. In contrast, Debater-4B continues to show incremental improvements when the length of CoD extends to 8, indicating that larger models benefit more from extended reasoning due to their stronger ability to integrate and retain detailed intermediate steps. Nonetheless, further increasing the CoD beyond a certain point (e.g., 12) may lead to overthinking and result in performance degradation for both model sizes. These observations demonstrate that while moderate depths effectively boost retrieval accuracy, excessively long chains can dilute the benefits and introduce unnecessary computational overhead. Overall, these findings underscore the importance of carefully tuning the thinking depth for LLM-based retrievers.
(a) TREC-COVID.
(b) FiQA.
(c) NFCorpus.
(d) SCIDOCS.
Figure 4: Performance of Debater at Different Thinking Steps. We collect all documents from each thinking step to demonstrate the retrieval performance of Debater across different stages of reasoning.
5.4 Retrieval Performance of CoD-Generated Document Representations
In this subsection, we investigate how the Chain-of-Deliberation (CoD) enhances the representation capability of LLM-based retrievers. Specifically, we evaluate the quality of embeddings produced at different thinking steps in CoD and assess their retrieval performance individually.
As shown in Figure 4, during the early stages of reasoning (e.g., steps 1 and 2), retrieval performance is relatively low and even decreases. This suggests that initial embeddings, based on minimal deliberation, may lack the nuanced understanding required for effective retrieval. However, as the number of thinking steps increases, performance generally improves, indicating that more deliberation leads to more refined embeddings for retrieval tasks. These improvements demonstrate that Debater helps LLMs better capture relevant information from the tokens of documents and previous decoded representations in the thinking steps. On the other hand, Debater’s performance gradually reaches a plateau as the thinking steps increase, suggesting that after a certain point, the embeddings become sufficiently fine-grained, and further deliberation may result in limited improvements and unnecessary computational complexity.
5.5 Characteristics of the Embeddings Generated by Debater
In this subsection, we analyze the embeddings learned by Debater from CoD. Specifically, we compute the average cosine similarity scores of embeddings generated at different positions in FiQA to understand how embeddings at various stages affect the final representation used for retrieval.
(a) Debater w/o SD.
(b) Debater.
Figure 5: Similarity Relationship Between Adjacent Position Embeddings. Both Debater w/o SD and Debater are compared in our experiments. Darker blue indicates a higher similarity score.
(a) Debater w/o SD.
(b) Debater.
Figure 6: Similarity Scores Between the First Seven Embeddings and the Last Embedding. The last embedding is used as the representation of documents for retrieval.
Learning Patterns of CoD. As shown in Figure 5(b), we present the average similarity scores among the first five embeddings generated by Debater to explore how Debater refines document representations step by step during CoD.
The results reveal a clear pattern in the similarity relationships: each embedding is most similar to its immediate neighbors, with similarity gradually decreasing as the distance between embeddings increases. This indicates that each embedding heavily relies on the previously decoded representations to generate more refined embeddings, which likely results from the autoregressive decoding mechanism of LLMs. Comparing this with the Debater w/o SD model (Figure 5(a)), we observe that the Debater model shows higher similarity scores with representations from more recent steps during CoD. This suggests that our Self Distillation method effectively encourages LLMs to learn more diverse representations at different thinking steps and to gather more relevant information from nearby steps, which leads to finer-grained document representations.
Contributions of CoD Steps to Document Representations. Figure 6 illustrates the similarity relationship between embeddings at intermediate thinking steps of CoD and the final document representation generated at the last thinking step. This helps us explore the contributions of different thinking steps to the final document representations.
In general, both Debater w/o SD and Debater models exhibit a trend of gradually increasing similarity to the final embedding as the thinking steps progress. As shown in Figure 6(a), the Debater w/o SD model tends to produce similar similarity scores with the final step, showing that relying solely on CoD may degrade the performance of Debater. It may lie in that all CoD generated embeddings are supervised with the same training loss and optimized to match the same query, making them become homogeneous. In contrast, Debater (Figure 6(b)) shows a more significant increase in similarity, indicating that these thinking steps contribute more variably to the final document representation. Notably, the information generated at each CoD step is gradually compressed into the last embedding, which further demonstrates the effectiveness of Debater in leveraging the thinking capacity of LLMs to generate more effective document representations for retrieval.
6 Conclusion
This paper proposed the Deliberate Thinking based Dense Retriever (Debater), a novel method designed to enhance the reasoning capabilities of LLM-based dense retrievers via deliberation-augmented embedding. Through the integration of Chain-of-Deliberation (CoD) and Self Distillation (SD), Debater significantly improves retrieval performance by capturing different views of documents before generating final embeddings. Our experimental results demonstrate that Debater outperforms existing dense retrievers by implementing with the LLM of a smaller scale.
Limitation
Debater demonstrates its effectiveness in activating the reasoning capabilities of LLMs to enhance document representation. However, each thinking embedding requires interaction with the query embedding during the training of Debater, which introduces additional memory costs for GPU. Furthermore, we only employ 2$\times$ NVIDIA A100-40G GPU to train and evaluate our Debater models. Due to the high embedding dimension of the LLM-based retriever, it takes us over a week to build the index for the BEIR evaluation using Debater-2.4B, and over 2 weeks on Debater-4B. Thus, we do not further explore larger Debater models (e.g., Debater-7B) in this paper due to limitations in computational resources. However, our experimental results indicate that the Debater-4B model outperforms all 7B-scale baseline models, highlighting its effectiveness.
Appendix A Appendix
A.1 Licenses
The E5 dataset and BEIR benchmark are released under the Apache License 2.0. The terms of use can be found on their Github pages. This license allows users to freely use, modify, and distribute the data, permitting academic use of their dataset.
A.2 More Details of Datasets Used in Experiments
In this section, we provide a detailed description of the training dataset and the evaluation benchmark for our Debater.
E5 Dataset. We employ the E5 dataset Wang et al. (2024a); Springer et al. (2025) for training, comprising a collection of publicly available datasets. The E5 dataset is carefully curated to encompass a wide range of retrieval scenarios and tasks. We present the statistics of the E5 dataset in Table 3. The instructions used in the E5 dataset are shown in Table 5. The full dataset can be available at their website [^2] https://github.com/jakespringer/echo-embeddings.
BEIR Benchmark. We evaluate our Debater models using the BEIR benchmark, which includes 18 datasets from 9 heterogeneous retrieval tasks. We focus on the 14 publicly available datasets and use the standard instructions in MTEB Muennighoff et al. (2023). The statistics for these datasets are provided in Table 4, and the instructions can be found in Table 6.
Dataset #Samples Proportion
ELI5 Fan et al. (2019) 32,547 2.16%
HotpotQA Yang et al. (2018) 90,447 5.99%
FEVER Thorne et al. (2018) 101,578 6.73%
MIRACL Zhang et al. (2023b) 32,561 2.16%
MSMARCO Passage Ranking Bajaj et al. (2016) 249,592 16.53%
MSMARCO Document Ranking Bajaj et al. (2016) 73,400 4.86%
NQ Kwiatkowski et al. (2019) 100,231 6.64%
NLI Gao et al. (2021) 277,230 18.36%
SQuAD Rajpurkar et al. (2016) 87,599 5.80%
TriviaQA Joshi et al. (2017) 73,346 4.86%
Quora Duplicate Questions DataCanary et al. (2017) 101,762 6.74%
Mr-TyDi Zhang et al. (2021) 48,715 3.23%
DuReader He et al. (2018) 86,395 5.72%
T2Ranking Xie et al. (2023) 154,294 10.22%
Table 3: Data Statistics of E5 Dataset. We show the composition and distribution of E5 Dataset.
Dataset Task Domain #Query #Corpus Avg. D / Q
TREC-COVID Bio-Medical Bio-Medical 50 171,332 493.5
NFCorpus Information Bio-Medical 323 3,633 38.2
NQ Question Wikipedia 3,452 2,681,468 1.2
HotpotQA Answering Wikipedia 7,405 5,233,329 2.0
FiQA-2018 (QA) Finance 648 57,638 2.6
ArguAna Argument Misc. 1,406 8,674 1.0
Touché-2020 Retrieval Misc. 49 382,545 19.0
CQADupStack Duplicate-Question StackEx. 13,145 457,199 1.4
Quora Retrieval Quora 10,000 522,931 1.6
DBPedia Entity-Retrieval Wikipedia 400 4,635,922 38.2
SCIDOCS Citation-Prediction Scientific 1,000 25,657 4.9
FEVER Fact Checking Wikipedia 6,666 5,416,568 1.2
Climate-FEVER Wikipedia Fact Checking 1,535 5,416,593 3.0
SciFact Scientific Fact Checking 300 5,183 1.1
Table 4: Data Statistics of the BEIR Benchmark.
Datasets Instructions
DuReader Given a Chinese search query, retrieve web passages that answer the question.
ELI5 Provided a user question, retrieve the highest voted answers on the Reddit ELI5 forum.
FEVER Given a claim, retrieve documents that support or refute the claim.
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question.
MIRACL Given a question, retrieve Wikipedia passages that answer the question.
MrTyDi Given a question, retrieve Wikipedia passages that answer the question.
MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query.
MSMARCO Document Given a web search query, retrieve relevant documents that answer the query.
NQ Given a question, retrieve Wikipedia passages that answer the question.
Squad Retrieve Wikipedia passages that answer the question.
T2Ranking Given a Chinese search query, retrieve web passages that answer the question.
TriviaQA Retrieve Wikipedia passages that answer the question.
QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question.
Find questions that have the same meaning as the input question.
NLI Given a premise, retrieve a hypothesis that is entailed by the premise.
Retrieve semantically similar text.
Table 5: Instructions Used for E5 Dataset.
Task Name Instruction
Arguana Given a claim, find documents that refute the claim
FiQA2018 Given a financial question, retrieve user replies that best answer the question
NFCorpus Given a question, retrieve relevant documents that best answer the question
Quora Given a question, retrieve questions that are semantically equivalent to the given question
SCIDOCS Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
TREC-COVID Given a query on COVID-19, retrieve documents that answer the query
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
ClimateFEVER Given a claim about climate change, retrieve documents that support or refute the claim
FEVER Given a claim, retrieve documents that support or refute the claim
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
DBPedia Given a query, retrieve relevant entity descriptions from DBPedia
NQ Given a question, retrieve Wikipedia passages that answer the question
MS MARCO Given a web search query, retrieve relevant passages that answer the query
CQADupStack Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
Table 6: Instructions Used for Evaluation on the BEIR Benchmark.
A.3 Case Study
Query In the US, is it a good idea to hire a tax consultant for doing taxes?
MiniCPM This may not exactly answer your question but, as a small business owner, I would highly recommend having a professional handle your taxes. It is worth the money to have it done correctly rather than doing something wrong and getting audited or worse having penalties assessed and owing more than you thought would be possible. I would recommend this especially if this is how you make your primary income, you can always write it off as a business expense.
Debater Whether you do decide to go with a tax advisor or not, be sure to do some research on your own. When we moved to the US about 5 years ago, I did find the taxes here pretty complicated and confusing ··· After all, they are also humans prone to mistakes and your taxes are your liability in the end. My suggestion is to start with a good tool that supports tax filing for non-residents. Most of them provide a step-by-step QA based tool. As you go through the steps, Google each question you don’t understand. It may take more time than hiring a tax advisor directly but in the end it will all be worth it.
Query What is the mechanism of inflammatory response and pathogenesis of COVID-19 cases?
MiniCPM The novel coronavirus disease (COVID-19) pandemic is placing significant strains on health systems, scientific communities, essential public services, and economies all over the world. In this context, the worlds scientific biomedical establishment is unleashing an unprecedented response to the COVID-19 pandemic. This is a battle against time, considering the thousands of human lives are lost every day. In this commentary, based on a very recent research report, we intend to highlight how a new mechanism describing the RAGE transactivation produced by Ang II-mediated ATR1 activation can run continuously and thus, reinforcing a sustained inflammation in lungs, due to the SARS-Cov-2-mediated imbalance of the ACE/And II/ATR1 pathway.
Debater The evidence on the pathophysiology of the novel coronavirus SARS-CoV-2 infection is rapidly growing. Understanding why some patients suffering from COVID-19 are getting so sick, while others are not, has become an informal imperative for researchers and clinicians around the globe. The answer to this question would allow rationalizing the fear surrounding this pandemic. Understanding of the pathophysiology of COVID-19 relies on an understanding of interplaying mechanisms, including SARS-CoV-2 virulence, human immune response, and complex inflammatory reactions with coagulation playing a major role. An interplay with bacterial co-infections, as well as the vascular system and microcirculation affected throughout the body should also be examined. More importantly, a comprehensive understanding of pathological mechanisms of COVID-19 will increase the efficacy of therapy and decrease mortality. Herewith, presented is the current state of knowledge on COVID-19: beginning from the virus, its transmission, and mechanisms of entry into the human body, through the pathological effects on the cellular level, up to immunological reaction, systemic and organ presentation. Last but not least, currently available and possible future therapeutic and diagnostic options are briefly commented on.
Table 7: Case Studies. We present two cases from FiQA and TREC-COVID, and show the top 1 passage retrieved from MiniCPM and Debater, with the key contents in the passage highlighted.
In this subsection, we present two case studies to demonstrate the effectiveness of Debater. Specifically, we show the top 1 retrieved document comes from both MiniCPM and Debater in the FiQA and Trec-COVID datasets. The retrieval cases are shown in Table 7.
For the query “In the US, is it a good idea to hire a tax consultant for doing taxes?”, while the retrieval result of MiniCPM suggests hiring a professional, it fails to mention the specific context of “In the US”. In contrast, the document retrieved by Debater aligns more closely with the query, addressing all the relevant aspects. This highlights that Debater effectively activates the reasoning capabilities of LLMs, enabling a more fine-grained document representation.
For the query “What is the mechanism of inflammatory response and pathogenesis of COVID-19 cases?”, MiniCPM retrieves documents containing more pathological terms, such as “RAGE transactivation” and “the ACE/Ang II/ATR1 pathway”. However, these retrieved documents cannot accurately address the inflammatory response and pathogenesis of COVID-19. This suggests that MiniCPM lacks a fine-grained understanding of document content, leading to suboptimal retrieval performance. In contrast, Debater provides accurate retrieval results, demonstrating the importance of deliberate thinking before searching.
A.4 Comparison with Additional Baseline Retrievers
In this subsection, we conduct additional comparisons with baseline retrievers not discussed in the main section, including several advanced retrievers from the BEIR benchmark.
Specifically, we use CPT Neelakantan et al. (2022), Udever Zhang et al. (2023a), ULLME, Ada2, and Promptriever as our additional retrievers to compare with Debater. CPT refers to a series of large encoder models developed by OpenAI Brown et al. (2020), which are pre-trained on large-scale and unsupervised data. We use the CPT-text-175B model as our baseline model. Udever contrastively trains the BLOOM Scao et al. (2022) to build dense retrievers, enabling it to effectively generate aligned multilingual embeddings for both text and code. ULLME Man et al. (2024) enables bidirectional attention between LLMs and enhances them for text embedding using Generative Reinforcement Learning (GRL). Ada2 is a versatile text embedding model introduced by OpenAI. The evaluation result for Ada2 comes from Kamalloo et al. (2023). Promptriever Weller et al. (2025) is trained on an instance-level instruction dataset from MS MARCO, enhancing its ability to follow instructions for retrieval tasks.
The performance comparison results on the BEIR benchmark are shown in Table 8. Compared with these baseline models, Debater still exhibits competitive performance, demonstrating its effectiveness. Compared to Udever and ULLME, Debater shows significant improvement on FEVER and Climate-FEVER, indicating its effectiveness across diverse fact-checking scenarios. When compared to Promptriever, Debater demonstrates similar retrieval performance. However, unlike Promptriever, Debater does not rely on synthetic data for training its retrieval models.
Method ($\rightarrow$) CPT Udever ULLME Ada2 Promptriever Debater
Model Size ($\rightarrow$) 175B 7B 7B unk. 7B 2.4B 4B
TREC-COVID 0.649 0.838 0.836 0.813 0.839 0.795 0.836
NFCorpus 0.407 0.360 0.394 0.358 0.365 0.378 0.399
NQ / 0.533 0.614 0.482 0.619 0.560 0.561
HotpotQA 0.688 0.567 0.674 0.654 0.692 0.678 0.678
FiQA 0.512 0.367 0.423 0.411 0.459 0.434 0.462
ArguAna 0.435 0.522 0.468 0.567 0.518 0.567 0.562
Touché-2020 0.291 0.252 0.271 0.280 0.314 0.211 0.250
Quora 0.638 0.883 0.878 0.876 0.865 0.886 0.886
DBPedia 0.408 0.376 0.464 0.402 0.450 0.430 0.432
SCIDOCS / 0.189 0.211 0.186 0.173 0.197 0.212
FEVER 0.775 0.740 0.615 0.773 0.828 0.859 0.857
Climate-FEVER 0.223 0.268 0.222 0.237 0.276 0.303 0.294
SciFact 0.754 0.693 0.724 0.736 0.750 0.735 0.743
CQADupStack / 0.393 / 0.391 / 0.431 0.428
Avg CPT sub 0.525 0.533 0.543 0.555 0.571 0.571 0.582
Avg / 0.499 / / / 0.533 0.543
Table 8: Retrieval Performances of Debater and Other Baseline Models on the BEIR Benchmark.